Let V= R^3

W={(x1, x2, x3)| x1-x2 =x3}. Show that W is a

subspace of V. Further,find a basis for W and hence,find the dimension of W.
1
Expert's answer
2021-08-25T13:32:39-0400

Let us show that "W=\{(x_1, x_2, x_3)| x_1-x_2 =x_3\}" is a subspace of "V". Let "(x_1, x_2, x_3),(y_1, y_2, y_3)\in W,\ a\in \R." Then "x_1-x_2=x_3,\ y_1-y_2=y_3." It follows that "a(x_1-x_2)=ax_3," and hence "ax_1-ax_2=ax_3." We conclude that "a(x_1, x_2, x_3)\in W." It follows also that "x_1-x_2+y_1-y_2=x_3+y_3," and thus "(x_1+y_1)-(x_2+y_2)=x_3+y_3." Therefore, "(x_1, x_2, x_3)+(y_1, y_2, y_3)\in W." We conclude that "W" is a subspace of "V."

Taking into account that

"W=\{(x_1, x_2, x_3)\in\R^3| x_1-x_2 =x_3\}=\{(x_2+x_3, x_2, x_3)| x_2, x_3\in\R\}\\
=\{x_2(1, 1, 0)+x_3(1, 0, 1)| x_2, x_3\in\R\},"

we conclude that "\{(1, 1, 0),(1, 0, 1)\}" is a basis of "W," and hence "dim(W)=2."